OpenStudy (anonymous):

use the function f(x)=2-5x to solve this.

OpenStudy (anonymous):

to solve?

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

oops what

OpenStudy (anonymous):

replace those x numbers in you chart for x in f(x) one at a time and solve for the corresponding y value I'll do the first to show you: 2- 5(-2) = 2 - (-10) = 2 + 10 =12 (-2, 12) is the ordered pair at x of -2 never mind i thought o made a mistake but I was mistaken

OpenStudy (anonymous):

ok well thank you so much!

OpenStudy (anonymous):

f(x) means to evaluate the function at x. and solving it gives you the y value at the x value examined

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

HELLO KAT